Operations Report – September 6th

MDRS Complete Operations Report 06-SEPT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Beginning of field season/work party daily update

Non-nominal systems: Nothing to report

Action is taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Operating

Solar— Nominal, providing all power
Solar— SOC % Last 24 hours:
Average 79.7
Minimum 62
Maximum 100
Note on solar: Nothing to report

Diesel Reading – 0 %.

Propane Reading, station tank – 80 %
Propane Reading, director tank— 75%
Propane Reading, intern tank— 72%
Propane Reading, generator: 80%

Ethanol Free Gasoline – not checked gallons

Water (loft tank): 50 gallons
Water Meter: n/a units
Water (static tank) – 150 gallons
Static to Loft Pump used – yes
Water in GreenHab – 175 gallons
Water in ScienceDome: 0 gallons
Water (Outpost tank) – 150 gallons

Hab toilet tank emptied: yes

Perseverance rover used: Yes
Beginning charge: 100
Ending charge:
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): All 350’s used
Reason for use: getting rocks for the tourist kiosk, geology
Oil Added? No
ATV Fuel Used: Not checked Gals
# Hours the ATVs were Used today: 0
Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, hauling water

CrewCar used and why, where? Yes, travel back and forth from town storage units

Luna used and why, where? No

General notes and comments: Nothing to report

Summary of internet: All nominal.

Summary of suits and radios: Spacesuits are offsite. Radios are charging.

Campus-wide inspection, if action is taken, what and why? Nothing to report

Summary of general operations: We built the pathway for the new tourist kiosk today. We also constructed part of the new tunnel coverings.

We did another purge of the storage units, this time going through and removing materials left over from the construction of the Hab! We also decided to remove things we had been storing for past crews, researchers, and expeditions. These are things that had been unclaimed for years.

I’m happy to report that much of this is finding new homes in Hanksville. We are still far from finished with the work there, but it’s a good start. The mud is dried except for where it has been trapped under or inside things but it smells really bad.

The goal is to remove everything from one unit, get what remakes all on shelves in a single unit, and then cleaned the floors of both. There will be room for storage of some things with the URC materials but I would like to get all things of use back at MDRS. For example, all of our power tools were in there but we could not access them.

Most of the damaged Ikea material was removed from packaging and cleaned and dried. We will be able to use most of it still.

Summary of Hab operations: Nothing to report

Summary of Outpost operations: Prep has begun for the new wall in the intern trailer bedroom. We have begun troubleshooting the intern trailer plumbing.

Summary of GreenHab operations: Water prep of the soil is underway. It’s still very hot to start seeding but I think it can be done. A few large herbs have been planted.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing at this time.

Copyright © The Mars Society. All rights reserved. | Main Site